(1 point) solve the system 18-21 dx X dt 14-17 with the initial value -16 x(0) = ༈ ༈་ -12 xt) =
Answer #1

- 306 +ith - 18n tu2294=0 Given that 18-21 n at 14 -17 sef Az 18 -21 14 -17 A The eigen value of A are det (A-n 12) =0 co 18now, the eigennector wenst the eigen value -3 is y (47353) (03)-18) 1963) -21 - 14 -14 21n-2ly=0 144-144 20 7 x (t) = ((!) **